OpenStudy (anonymous):

(2x)/(x^2-x-2) - (4x)/(x^2-3x+2) simplify the difference...

OpenStudy (anonymous):

Solve each side first....

OpenStudy (pfenn1):

\[\frac{2x}{x^2-x-2} - \frac{4x}{x^2-3x+2}\] The first thing I would do is factor the denominators.

OpenStudy (anonymous):

so it would be \[2x \div(x+1)(x-2) - 4x \div(x-1)(x-2)\] right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

okay and then i have to find the common denominator?

OpenStudy (pfenn1):

Yes

OpenStudy (anonymous):

Yes Ma'am

OpenStudy (anonymous):

and that would be x-2?

OpenStudy (anonymous):

Yep! (:

OpenStudy (anonymous):

so then i multiply each side with x-2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so then it would be (x-2)/(x-2) * (2x)/(x+1)(x-2) for the first side. right? does any of the (x-2)'s cancel out?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No, because it isn't in the correct form for them to cancel out (:

OpenStudy (anonymous):

okay so then it would be \[2x ^{2}-4x/x ^{3}-3x ^{2}+4\] on the first side right?

OpenStudy (anonymous):

and then on the second side it would be \[4x ^{2}-8x/x ^{3}-5x ^{2}+8x-4\] right?

OpenStudy (anonymous):

Correct !! (:

OpenStudy (anonymous):

Okay. I don't know what to do after that.

OpenStudy (anonymous):

Where exactly are u at with this problem?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Write out everything you've answered so far..

OpenStudy (anonymous):

okay i've got \[2x/(x ^{2}-x-2) - 4x/(x ^{2}-3x+2)\] as the problem. then i multiplied (x-2) to each side after I factored the denominators. so it looked like this: \[(x-2)/(x-2) * 2x/(x+1)(x-2) - 4x/(x-1)(x+2) * (x-2)/(x-2)\] and after i multiplied them i got \[(2x ^{2}-4x)/(x ^{3}-3x+4) - (4x ^{2}-8x)/(x ^{3}-5x ^{2}+8x-4)\]

OpenStudy (anonymous):

and thats all i have so far

OpenStudy (anonymous):

Shucks, I'm lost!

OpenStudy (anonymous):

Haha, me too.

OpenStudy (asnaseer):

You worked correctly up to this point:\[\frac{2x}{(x+1)(x−2)}−\frac{4x}{(x−1)(x−2)}\]

OpenStudy (asnaseer):

after this, you correctly noticed the common factor between the two denominators was (x-2).

OpenStudy (asnaseer):

what you missed was calculating the LCM of the denominators as the next step.

OpenStudy (asnaseer):

so we have the denominators as:\[(x+1)(x-2)=(x+1)\times(x-2)\]and:\[(x-1)(x-2)=(x-1)\times(x-2)\]so the LCM would be:\[(x+1)(x-1)(x-2)\]

OpenStudy (asnaseer):

a strategy you can use to find the LCM is as follows: 1. remove all common factor(s) - in this example it was (x+2) 2. multiply all remaining terms together with the removed common factor(s) - in this example it was (x+1)(x-1) multiplied by the common factor of (x+2)

OpenStudy (anonymous):

okay so now i multiply each side by all three of those?

OpenStudy (asnaseer):

no, the next step is to identify what you need to multiply the denominator by in order to get the LCM. so, for the first term, the denominator is (x+1)(x-2) and the LCM is (x+1)(x-1)(x-2), which means we need to multiply the first fraction by (x-1)/(x-1)

OpenStudy (asnaseer):

so first fraction becomes:\[\frac{2x}{(x+1)(x−2)}*\frac{x-1}{x-1}=\frac{2x(x-1)}{(x+1)(x-1)(x−2)}\]

OpenStudy (asnaseer):

now what do you think we need to multiply the second fraction by?

OpenStudy (asnaseer):

remember the second denominator is (x-1)(x-2) and we know the LCM is (x+1)(x-1)(x-2). so what do we need to multiply (x-1)(x-2) by in order to get (x+1)(x-1)(x-2)?

OpenStudy (anonymous):

(x+1)?

OpenStudy (asnaseer):

correct, so the second fraction becomes:\[\frac{4x}{(x−1)(x−2)}=\frac{4x}{(x−1)(x−2)}*\frac{x+1}{x+1}=\frac{4x(x+1)}{(x+1)(x−1)(x−2)}\]

OpenStudy (asnaseer):

do you think you can complete the question from here now?

OpenStudy (anonymous):

i think i can try.

OpenStudy (asnaseer):

good - have a go at it and post your working here - I'll come back and check it for you later - good luck! :)

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

so would the answer be \[-2x ^{2}+2x/(x+1)(x-1)(x-2)\] right?

OpenStudy (asnaseer):

not quite, it should be:\[\frac{2x(x-1)}{(x+1)(x-1)(x−2)}-\frac{4x(x+1)}{(x+1)(x−1)(x−2)}=\frac{2x(x-1)-4x(x+1)}{(x+1)(x-1)(x−2)}\]\[\qquad=\frac{2x^2-2x-4x^2-4x}{(x+1)(x-1)(x−2)}\]\[\qquad=\frac{-2x^2-6x}{(x+1)(x-1)(x−2)}\]\[\qquad=\frac{-2x(x+3)}{(x+1)(x-1)(x−2)}\]

OpenStudy (anonymous):

alright. I see i forgot the negative on one thing. Well thanks very much! (:

OpenStudy (asnaseer):

yw